Kevin Connell, Timothy Parker, Nicholas Zerlentes

1. Who first came up with the Four Color Theorem?

1. What is the historical significance of this theorem?

1. What is the misconception of this theorem? Marissa Eyre, Joe Fernandez, and Tracey Sabey’s project ’Peg Solitaire.’

1. In the strategy described with pegs as 1 and an empty slot as 0, what is a possible position coming before 011?

2. Why does one have to end in a y-position to win the game?

3. Using the method we described, how many possible places can the last peg land to win?